1. A UHF television loop antenna has a diameter of 11 cm. The magnetic field of a TV signal is normal to the plane of the loop and, at one instant of time, its magnitude is changing at the rate 0.16 T/s. The magnetic field is uniform. What emf is induced in the antenna?



1
Expert's answer
2022-03-23T17:59:26-0400

The EMF induced in the antena

"{\cal E}=\frac{d\Phi}{dt}=\frac{dB}{dt}A=\frac{dB}{dt}\pi d^2/4\\
=0.16*3.14*0.11^2/4=1.5*10^{-3}\:\rm V"